Circular No. 6038

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope (HST) Team, K. S. Noll, Space Tele-

scope Science Institute, reports the following results: “In a spectrum of the G impact site obtained on July 21.67 UT, we observe strong emission from Mg^{+} at 279.9 and 280.4 nm, from neutral Mg at 285.2 nm, and from CS at 257.8 nm. At least six weaker emission lines correspond to neutral iron. Additional weak spectral features may correspond to emission from other atoms or molecules. The G impact site was close to the limb at the time of our observations, and a portion of our aperture may have extended beyond Jupiter’s limb.”

R. Hunstead and D. Campbell-Wilson, University of Sydney, report the results of observations of Jupiter over the period July 12–13 and 15–23 using the Molonglo Observatory Synthesis Telescope (MOST) at 843 MHz: “The planet was tracked for 9 hr (4.5 hr on either side of ∼ 8^{h}30^{m} UT each day). Preliminary images have been formed and averaged over each 9-hr integration. Integrated flux densities have been measured in a consistent way on each image and show an essentially constant value over the period July 12–16, followed by a steady increase in flux density over the 7-day duration of the SL9 impacts. The overall increase of 50 percent from July 16 to 23 is highly significant and completely unexpected. Further monitoring is continuing with MOST, and a more detailed analysis of these (and earlier) observations is underway. Other observatories, particularly those operating at frequencies < 1 GHz, are encouraged to continue monitoring Jupiter to define better the spectrum of the enhanced emission.”

PSR 0833–45 C. Flanagan, Hartebeesthoek Radio Astronomy Observatory, Johan-

nesburg, reports: “PSR 0833–45 (the Vela pulsar) underwent a spin-up of relative magnitude dP/P = −8.35(2) × 10^{−7} on July 26.057 UT. This is the tenth such large ‘glitch’ observed from this pulsar. Observations at Hartebeesthoek commenced 4.3 hr after the extrapolated glitch epoch.”

P. M. McCulloch, University of Tasmania, writes: “The frequency of the Vela pulsar PSR 0833–45 increased abruptly around July 26.060 UT; a change in pulse frequency of 9.635 × 10^{−6} Hz occurred at this time. The glitch, which is of similar magnitude to previous events, was observed at 635 and 990 MHz from the Mount Pleasant radio astronomy observatory, Hobart.”

1994 July 27 Daniel W. E. Green